INFRARED SPECTROSCOPY OF PROTONATED CARBONYLS: PROTONATED GLYOXAL

\underline{J. D. MOSLEY}, T. C. CHENG, and M. A. DUNCAN, University of Georgia, Dept. of Chemistry, 1001 Cedar St, Athens, GA 30602.

The protonation site of carbonyls is usually at oxygen, allowing studies of the O − H^{+} and carbonyl stretch vibrations and their variation with the local chemical environment. We have already studied protonated acetone and its protonbridged dimer and now extend the study of protonated carbonyls to protonated glyoxal. Glyoxal is the simplest αoxoaldehyde, which has the chance to form an intramolecular bridging proton structure upon protonation. Computational chemistry predicts the proton-bridged cis isomer to be the lowest energy structure. The infrared spectrum of mass-selected protonated glyoxal is obtained using infrared photodissociation spectroscopy with Ar tagging. The spectrum shows the presence of only the higher energy trans isomer when compared with purely harmonic calculations. The frequencies of the O − H^{+} and carbonyl stretch vibrations of protonated glyoxal are compared to other protonated carbonyls.